# Echo: Decoupling Inference and Training for Large-Scale RL Alignment on Heterogeneous Swarms

### v1

@@ Abstract
Modern RL-based post-training for large language models (LLMs) co-locate trajectory sampling and policy optimisation on the same GPU cluster, forcing the system to switch between inference and training workloads. This serial context switching violates the single-program–multiple-data (SPMD) assumption underlying today’s distributed training systems. We present Echo, the RL system that cleanly decouples these two phases across heterogeneous “inference” and “training” swarms while preserving statistical efficiency. Echo introduces two lightweight synchronization protocols: a sequential pull mode that refreshes sampler weights on every API call for minimal bias, and an asynchronous push–pull mode that streams version-tagged rollouts through a replay buffer to maximise hardware utilisation. Training three representative RL workloads with Qwen3-4B, Qwen2.5-7B, and Qwen3-32B on a geographically distributed cluster, Echo matches a fully co-located Verl baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware. These promising results demonstrate that large-scale RL for LLMs could achieve datacentre-grade performance using decentralised, heterogeneous resources.
@@ 1 Introduction
Reinforcement learning (RL) has become the de facto recipe for aligning billion-parameter language models with human preferences, such as RLHFouyang2022training , RLAIFlee2023rlaif , DPOrafailov2023direct and their rapidly proliferating variants. All of these algorithms share a multi-program, multi-data (MPMD) execution pattern: (i) an inference phase that generates trajectories under the current policy, followed by (ii) a training phase that converts those trajectories into gradient updates. State-of-the-art systems, e.g., DeepSpeed RLHFyao2023deepspeed and Verlsheng2024hybridflow , run the two phases on the same high-bandwidth cluster, implicitly synchronising through global barriers and expensive NVLink / InfiniBand fabrics. Since trajectory sampling and policy optimization have very different hardware footprints, running both phases on a single hardware pool forces the cluster to serially contextual-switch between two conflicting workloads. This back-and-forth violates the single-program–multiple-data (SPMD) assumption that underpins most large-scale GPU stacks, hindering utilisation and scalability.
@@
In this paper, we introduce Echo, the RL system that solves the above problem through a pair of synchronization mechanisms, as shown in Figure 1. The sequential mode has trainers pull rollouts via an API, refreshing sampler weights beforehand to mirror classic PPO and minimise bias. The asynchronous mode lets samplers push tagged rollouts into a replay buffer while a coordinator caps version drift, maximising device utilisation.
Besides, on the inference side we build upon Parallax, a fully decentralised pipeline-parallel engine that converts geographically dispersed consumer GPUs and Apple-Silicon devices into a single, high-throughput sampler via dynamic KV-cache management and continuous batching. On the training side we adopt the community-standard Verl stack, adding robust support for LoRA-style parameter-efficient finetuning.
@@ 2.1 System Overview
Conventional RL stacks co-locate trajectory generation and policy updates on the same high-bandwidth cluster, implicitly relying on barrier-style synchronisation to guarantee that every gradient step consumes up-to-date data. Echo deliberately splits these two phases onto distinct, heterogeneous swarms—an inference swarm that samples trajectories and a training swarm that optimises the policy—so that each can run on the hardware best suited to its workload (edge devices versus datacenter GPUs). The price of this flexibility is a non-trivial information synchronisation problem: trajectory freshness and weight staleness must be controlled explicitly; otherwise, the decoupled system may diverge or converge more slowly than monolithic baselines. Tackling this challenge requires a careful analysis of the data-dependency chain between policy parameters, sampled trajectories, and gradient updates.

### v2

@@ Abstract
Modern RL-based post-training for large language models (LLMs) co-locate trajectory sampling and policy optimisation on the same GPU cluster, forcing the system to switch between inference and training workloads. This serial context switching violates the single-program–multiple-data (SPMD) assumption underlying today’s distributed training systems. We present Echo, the RL system that cleanly decouples these two phases across heterogeneous “inference” and “training” swarms while preserving statistical efficiency. Echo introduces two lightweight synchronization protocols: a sequential pull mode that refreshes policy weights according to API call for minimal bias, and an asynchronous push–pull mode that streams version-tagged rollouts through a replay buffer to maximise hardware utilisation. Training three representative RL workloads with Qwen3-4B, Qwen2.5-7B, and Qwen3-32B on a geographically distributed cluster, Echo matches a fully co-located Verl baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware. These promising results demonstrate that large-scale RL for LLMs could achieve datacentre-grade performance using decentralised, heterogeneous resources.
[…]
@@ 1 Introduction
Reinforcement learning (RL) has become the de facto recipe for aligning billion-parameter language models with human preferences, such as RLHFouyang2022training , RLAIFlee2023rlaif , DPOrafailov2023direct and their variants. All of these algorithms share a multi-program, multi-data (MPMD) execution pattern: (i) an inference phase that generates trajectories under the current policy, followed by (ii) a training phase that converts those trajectories into gradient updates. State-of-the-art systems, e.g., DeepSpeed RLHFyao2023deepspeed and Verlsheng2024hybridflow , run the two phases on the same high-bandwidth cluster, implicitly synchronising through global barriers and expensive NVLink / InfiniBand fabrics. Since trajectory sampling and policy optimization have very different hardware footprints, running both phases on a single hardware pool forces the cluster to serially contextual-switch between two different workloads. This back-and-forth violates the single-program–multiple-data (SPMD) assumption that underpins most large-scale training systems, hindering utilisation and scalability.
@@
In this paper, we introduce Echo, the RL system that solves the above problem through a pair of synchronization mechanisms, as shown in Figure 1. The sequential mode has trainers pull rollouts via an API, refreshing policy weights beforehand to mirror classic PPO and minimise bias. The asynchronous mode lets samplers push tagged rollouts into a replay buffer while a coordinator keeps tracking version drift, maximising device utilisation.
On the inference side, we build upon Parallax, a fully decentralised pipeline-parallel engine developped by Gradient Network that converts geographically dispersed consumer GPUs and Apple-Silicon devices into a single, high-throughput sampler via dynamic KV-cache management and continuous batching. On the training side we adopt the community-standard Verl stack, adding robust support for LoRA-style parameter-efficient finetuning.
@@ 2.1 System Overview
Conventional RL stacks co-locate trajectory generation and policy updates on the same high-bandwidth cluster, implicitly relying on barrier-style synchronisation to guarantee that every gradient step consumes up-to-date data. Echo deliberately splits these two phases onto distinct, heterogeneous swarms—an inference swarm that samples trajectories and a training swarm that optimises the policy—so that each can run on the hardware best suited to its workload. The price of this flexibility is a non-trivial information synchronisation problem: trajectory freshness and weight staleness must be controlled explicitly; otherwise, the decoupled system may diverge or converge more slowly than monolithic baselines. Tackling this challenge requires a careful analysis of the data-dependency chain between policy parameters, sampled trajectories, and gradient updates.